OpenStudy (anonymous):

K & m are supplementary angles . If m < K=3x+8 and m

OpenStudy (anonymous):

supplementary angles add up to 180 degrees so set them equal to 180... 3x + 8 + x + 12 = 180 -- combine like terms 4x + 20 = 180 -- subtract 20 4x = 160 -- divide by 4 x = 160/4 x = 40 m < K = 3x + 8 3(40) + 8 120 + 8 128 m < M = x + 12 40 + 12 52 check.. 128 + 52 = 180 180 = 180 (correct) < K = 128 and < M = 52